NvidiaIntely (nvidiaintely):

On a vacation, DeVante and his family traveled 385 miles at 65 mph. How long did the trip take? 5.96 hours 5.59 hours 5.69 hours 5.92 hours

OpenStudy (kevin):

v = s/t 65 = 385/t t = 5.92

NvidiaIntely (nvidiaintely):

Thanks!